OpenStudy (anonymous):

a woman collects $2.75 for each burrito sold but pays 25 dollars a day for a booth. how much does she leave with at the end of the day? ( assuming she came with nothing). write a function for it if it can be done

OpenStudy (anonymous):

2.75x-25=x

OpenStudy (anonymous):

Thank you, can you help with 4 more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay heres the next one. a car rental agency will rent you a subcompact for 27.50 a day and the per mile charge is 10 cents. what would be the function of the cost? same thing for all write the function if it can be written

OpenStudy (anonymous):

27.50x+10x=x ? i think

OpenStudy (anonymous):

okay and the next one is there are 3 points on a graph that are lined up. they are (0,10), (1,20), (2,30). write a function for that

OpenStudy (anonymous):

im not sure how to write that in a function one sec

OpenStudy (anonymous):

alright! well heres another one: netflix charges 8dollars per month and 2.99 per dvd mailed out. write a function for the total monthly cost.

OpenStudy (anonymous):

8+2.99x=x

OpenStudy (anonymous):

and the last one is a car rents for 30 dollars each day and 20 cents a mile. write a function for the total daily cost.

OpenStudy (anonymous):

30+20x=x

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

no problem!(: